# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ 4.3 Separation in Approximate Sampling
We first analyze the expressiveness of CoT. While algorithms such as FPAUS (and FPRAS) regard the accuracy parameter $\varepsilon$ as part of the input, this contrasts with the behavior of practical LLMs: once a base model has been pretrained to sufficiently low error, it can meet the required accuracy $\varepsilon$ at inference time. To formalize this, we first formally define well-pretrained models.
Definition 4.5 (Pretraining convergence).
Let $p$ be a target distribution supported on $\Sigma^{*}$, let $\alpha(n)\in\mathsf{poly}(n)$, and let $\gamma\in(0,\tfrac{1}{2})$. A stochastic model $\pi_{\omega}$, indexed by the internal randomness $\omega$ for each inference run, is said to be an $(\alpha,\gamma)$-weak probable approximation to $p$ if, for every position $i\in\{1,\ldots,m\}$, every input $x\in\Sigma^{*}$, and every prefix $y_{<i}\coloneq(y_{1},\ldots,y_{i-1})$, it holds that
$\Pr_{\omega}\left[(1-\tfrac{1}{\alpha(|x|)})p(y_{i}\mid x,y_{<i})\;\leq\;\pi_{\omega}(y_{i}\mid x,y_{<i})\;\leq\;(1+\tfrac{1}{\alpha(|x|)})p(y_{i}\mid x,y_{<i})\right]\;\geq\;\tfrac{1}{2}+\gamma.$
Intuitively, the model does not guarantee a small approximation error in every run, but achieves with probability strictly greater than random chance. We begin by showing the expressiveness of a pretrained CoT, augmented with inference-time algorithms, such as majority voting.
Proposition 4.6.
Let $R$ be a relation. Suppose the CoT model $\pi$ is an $(\alpha,\gamma)$-weak probable approximation to the uniform distribution $U(R(x))$ for a sufficiently large $\alpha$. Then there exists a polynomial-time algorithm, with access to $\pi$, that constitutes an FPAUS.
We now establish a conditional separation between CoT and Looped TFs. Building on the result of Nowak et al. (2024), we show that CoT can represent any FPRAS, and therefore can realize convergence during pretraining, and then constitute an FPAUS as above. In contrast, we show that there exist instances where a Looped TF fails to achieve comparable approximation guarantees.
@@
Assume $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations $R$, with respect to the counting problem $|R(x)|$. Then there exists a function $\alpha(n)\in\mathsf{poly}(n)$ and constant $\gamma>0$ such that a CoT model with polynomial steps satisfies the $(\alpha,\gamma)$-weak probable approximation to the uniform distribution $U(R(x))$, whereas no Looped TF with polynomial loops can achieve the same approximation, nor realize an FPAUS for $U(R(x))$.
@@
Suppose, toward a contradiction, that for every polynomial function, there exists a Looped TF satisfying the weak probable approximation. Since a Looped TF introduces randomness only at the decoding stage, its logits are computed deterministically, and the stochasticity can thus be derandomized. This would yield an FPTAS, contradicting the assumption that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$. ∎
Remark.
Formal proofs are given in Appendix D. This result suggests that looped models may not achieve the same expressive power as CoT in approximation under the assumption, uncovering an advantage of CoT and the benefit of stochastic decoding for enabling randomized approximation.
@@ 5 Experiments
In this section, we experimentally validate our theoretical results. For parallelizable tasks, Looped TF achieves more efficient solutions than CoT. In contrast, for computationally hard tasks, CoT leverages stochastic sampling to obtain approximate solutions, whereas Looped TF cannot achieve the same.
@@ 5.1 Experimental Setting
This section provides the task descriptions and training methodology for both tasks.
5.1.1 Parallelizable Tasks
Task Description
@@ 5.1.1 Parallelizable Tasks
We use four problems. (1) Word problems for finite non-solvable groups: given a sequence of generators, the task is to evaluate their composition. This problem is $\mathsf{NC}^{1}$-complete (Barrington, 1986), also studied for Looped TF by Merrill & Sabharwal (2025a). (2) sstt connectivity (STCON): given a directed graph $G=(V,E)$ and two vertices $s,t\in V$, the task is to decide whether $t$ is reachable from $s$. This problem belongs to $\mathsf{TC}^{1}$ (Gibbons & Rytter, 1989). (3) Arithmetic expression evaluation: given a formula consisting of $+,\times,-,/$ operations on integers, the task is to evaluate it. This problem is $\mathsf{TC}^{0}$-reducible to Boolean formula evaluation (Feng et al., 2023), which is $\mathsf{NC}^{1}$-complete (Buss, 1987). (4) Edit distance: given two strings $x$ and $y$, the task is to compute the minimum cost to transform $x$ into $y$. By reducing the dynamic programming formulation to shortest paths, this problem is in $\mathsf{TC}^{1}$ (Apostolico et al., 1990). Details are in Section E.1.1.
Training Methodology
For both CoT and Looped, the same Transformer block is adopted to ensure comparable parameter counts. For CoT models, training is performed with supervision of sequential algorithms. For instance, in dynamic programming tasks, the model is supervised to emit DP table entries. For inference, only the input is given. To control the number of CoT steps, intermediate positions are uniformly sampled, while preserving structure, along the full solution trajectory. All algorithms for each task are provided in Section E.1.2. We found that, just as CoT benefits from intermediate supervision, Looped similarly benefits from curriculum learning for certain tasks, as also discussed by Merrill & Sabharwal (2025a). Details of configuration are in Section E.1.3.
5.1.2 Approximation Task
Task Description
Since approximate counting and approximate sampling are equivalent for self-reducible relations, we employ approximate counting tasks for experimental evaluation, specifically the DNF counting task. The input is a propositional formula $F$, expressed as the disjunction of $m$ clauses $C_{1},\ldots,C_{m}$, where each clause is a conjunction of literals over $n$ Boolean variables. The output is the number of truth assignments over the $n$ variables that satisfy $F$, denoted by $|F|$. The DNF counting problem is $\mathrm{P}-$complete; thus, if $\mathsf{P}\neq\mathsf{NP}$, no polynomial-time algorithm exists for exact computation. However, a FPRAS has been established (Karp et al., 1989). In our experiments, we use $n=5$ variables, $m=10$ clauses, each of width $3$.
Training Methodology
For Looped TF, we trained the models end-to-end to directly map inputs to counts. In contrast, CoT models were trained with intermediate steps simulating a single trial of the randomized counting algorithm of Karp & Luby (1983). Performance is evaluated using relative error. At inference time, CoT executes multiple Monte Carlo samples and aggregates the results through post-processing. The total number of CoT steps is defined as the product of the reasoning steps per sample and the number of Monte Carlo samples. Further details are provided in Section E.2.
@@ 5.2 Results
Table 2: Accuracy on parallelizable tasks, with varying loop/step counts. Looped TF solves the tasks with fewer loops, whereas CoT requires more steps. Here, $n$ denotes the input size. $n=32/16$ indicates that Looped TF uses $n=32$ and CoT uses $n=16$.
Task $\bm{n}$ Looped TF (loops) CoT (steps)
2 4 6 8 8 16 32 64
Word Problem 64 0.8 0.8 100.0 100.0 0.8 0.8 0.8 100.0
Graph Connectivity 32 80.8 95.8 99.0 99.0 81.0 81.4 83.6 100.0
Arithmetic Evaluation 32/16 43.7 99.4 99.5 99.7 41.0 41.5 41.2 82.5
Edit Distance 32/16 57.3 72.9 86.2 90.7 69.2 70.3 82.6 94.8
Table 2 presents results on parallelizable tasks, comparing Looped TF and CoT under varying loop or step counts. Across all tasks, Looped TF solves the problems with only a small number of loops, whereas CoT requires substantially more steps to reach the same performance. These results support our theoretical results on the separation between Looped TF and CoT for parallel computation.
Figure 2: Accuracy of $r$-looped TFs on the arithmetic evaluation (left) and the connectivity (right). Each curve shows the performance for a fixed loop count $r$ as the input size $n$ increases. Increasing the loop count improves accuracy, and a logarithmic number of loops suffices.
We also follow the evaluation protocol of measuring performance across varying input sizes and loop counts, a methodology also adopted in prior studies on word problems (Merrill & Sabharwal, 2025a) and $k$-hop tasks (Sanford et al., 2024b). Figure 2 presents our results for Looped TFs, illustrating that as the input size $n$ increases, the number of loops required to maintain high accuracy grows only logarithmically. This empirical trend aligns with our theoretical claim that a (poly-)logarithmic loop count suffices, thereby demonstrating that Looped TFs can efficiently simulate parallel computation.
These results extend prior work in several directions. For the word problem, Merrill & Sabharwal (2025a) examined looped and standard Transformers, but our results show that a small number of CoT steps is insufficient to solve it. For graph tasks of Sanford et al. (2024b), we evaluate larger input sizes and show that looped models remain sufficient. For arithmetic evaluation and edit distance, while Feng et al. (2023) demonstrated that CoT suffices with quadratic steps, our results show that Looped TF can solve them more efficiently.
Figure 3: Relative error on DNF counting task.
Figure 3 shows the result on the approximation task. The CoT model achieves approximate inference via Monte Carlo sampling, and its relative error decreases markedly when provided with sufficient reasoning steps and samples. In contrast, the Looped TF fails to learn even as the number of loops increases, with the error remaining constant because the model outputs identical predictions across all inputs. These findings support our theoretical claim on the superiority of CoT in randomized approximation.
@@ 6 Conclusion
We have conducted a formal comparison between Chain-of-Thought (CoT) and Latent Thought in Looped Transformers, elucidating their respective strengths: Latent Thought enables efficient parallel computation, whereas CoT enables randomized approximation. This study characterizes these distinct computational paradigms of Transformer architectures from the perspective of representational capacity. These findings are further substantiated by experimental evidence. Although this study compared loop counts and step counts under equivalent settings, it did not explicitly address computational cost. An important direction for future work is to examine inference efficiency on hardware and to assess whether practical acceleration methods, such as parallelization, speculative decoding, and distillation, can be applied to reduce the required number of loops or steps.
Appendix A Notation
@@
Appendix B Formal Definitions for Section 2
B.1 Transformer Block
We define the computational components of a Transformer using the notation of Merrill & Sabharwal (2025a). Let $\mathbb{F}_{s}$ denote the set of $s$-bit fixed-point numbers with truncated arithmetic.
Definition B.1 (Transformer).
@@ A Transformer consists of the following components:
    A saturated multi-head self-attention layer $\mathbf{SA}:\mathbb{F}_{s}^{m\times N}\to\mathbb{F}_{s}^{m\times N}$ for arbitrary sequence length $N$, parameterized by a matrix ${\bm{O}}:\mathbb{F}_{s}^{s\times H}\to\mathbb{F}_{s}^{m}$ and, for each head $h\in[H]$ with head size $s$, matrices ${\bm{Q}}_{h},{\bm{K}}_{h},{\bm{V}}_{h}:\mathbb{F}_{s}^{m}\to\mathbb{F}_{s}^{s}$. Given an input ${\bm{x}}_{i}\in\mathbb{F}_{s}^{m}$ for each position $i\in[N]$, it computes the query $\mathbf{q}_{i,h}={\bm{Q}}_{h}{\bm{x}}_{i}$, key ${\bm{k}}_{i,h}={\bm{K}}_{h}{\bm{x}}_{i}$, and value $\mathbf{v}_{i,h}={\bm{V}}_{h}{\bm{x}}_{i}$, and outputs ${\bm{O}}\cdot({\bm{a}}_{i,1},\ldots,{\bm{a}}_{i,H}),$ where each attention output ${\bm{a}}_{i,h}$ is defined as:
    ${\bm{a}}_{i,h}=\sum_{j=1}^{c(i)}\frac{\exp(\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j,h})}{Z_{i,h}}\cdot\mathbf{v}_{j,h},\quad Z_{i,h}=\sum_{j=1}^{c(i)}\exp(\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j,h}),$
    with $c(i)=i$ for causal attention and $c(i)=n$ for full attention.
@@
B.2 Chain-of-Thought
Definition B.2 (CoT).
@@ B.2 Chain-of-Thought
$f_{\mathrm{dec}}\coloneq\mathbf{OUT}\circ(\operatorname{id}+\mathbf{FF}_{L})\circ(\operatorname{id}+\mathbf{SA}_{L})\circ\cdots\circ(\operatorname{id}+\mathbf{FF}_{1})\circ(\operatorname{id}+\mathbf{SA}_{1})\circ(\mathbf{WE}+\mathbf{PE}),$
@@
$f_{\mathrm{cot}}^{k+1}(x)\coloneq f_{\mathrm{cot}}^{k}(x)\cdot\mathrm{Dec}(f_{\mathrm{dec}}(f_{\mathrm{cot}}^{k}(x))),$
where $\cdot$ denotes concatenation of sequences, and $\mathrm{Dec}(\cdot)$ is a decoding function that maps the output logits to a token in $\mathcal{V}$}: in the deterministic model, $\mathrm{Dec}(z)\coloneq\arg\max_{i\in[|{\mathcal{V}}|]}z_{i}$; in the stochastic model, $\mathrm{Dec}(z)\sim\mathrm{Multinomial}({z_{i}}/{\sum_{j}z_{j}})$, assuming $z_{i}>0$ for all $i$. The final output of the CoT model after $T(n)$ steps is defined as the last output length $m$ tokens of $f_{\mathrm{cot}}^{T(n)}(x)$.
B.3 Looped Transformer
Definition B.3 (Looped TF).

### v2

@@ 4.3 Separation in Approximate Sampling
We then move from token-level conditional distributions $p(y_{i}\mid x,y_{<i})$ to the full sequence-level distribution $p(y\mid x)$. Beyond approximate counting, we establish a separation for approximate sampling problems. Specifically, we construct target distributions for which the complexity of each conditional can be reduced to approximate counting.
@@
Assume that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations. There exists a distribution $p(y\mid x)$ over $y\in\Sigma^{*}$ and $x\in\Sigma^{n}$ such that a CoT with a polynomial number of steps, whose induced output conditionals are denoted by $q(y_{i}\mid x,y_{<i})$, admits an FPRAS for approximating the conditional probabilities $p(y_{i}\mid x,y_{<i})$ for all $x\in\Sigma^{n}$, indices $i\geq 1$, and prefixes $y_{<i}\coloneq(y_{1},\ldots,y_{i-1})$. In contrast, no latent thought with polynomially many iterations admits the same approximation guarantee.
@@
Define the target distribution $p$ to be the uniform distribution supported on the solution set $R(x)$. We rely on the classical result that approximate sampling from the uniform distribution over solutions, captured by the class $\mathsf{FPAUS}$, is polynomial-time inter-reducible with approximate counting for self-reducible relations (Jerrum et al., 1986). Let $U(\cdot\mid x)$ denote the uniform distribution over solutions of a self-reducible relation $R(x)$. This distribution admits an autoregressive factorization $U(y\mid x)\;=\;\prod_{i=1}^{m}p(y_{i}\mid x,y_{<i}),$ where each conditional probability is given by $p(y_{i}\mid x,y_{<i})=\frac{\bigl|\{\,z\in\Sigma^{*}:(x,y_{1:i+1}z)\in R\,\}\bigr|}{\bigl|\{\,z\in\Sigma^{*}:(x,y_{1:i}z)\in R\,\}\bigr|}.$ We show that each conditional probability, expressed as a ratio of subproblem counts, reduces to approximate counting. Then, applying Lemma 4.3 to these conditionals yields the desired separation for approximate sampling. ∎
Figure 5: The separation for approximate counting (sampling).
Consequently, we obtain the following separations in favor of CoT, as also shown in Fig. 5.
Theorem 4.5.
Assuming $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations, it holds that
$\forall\,\mathcal{M}\in\{\mathsf{pCT},\mathsf{pLoop}\},\quad\mathcal{M}[\mathsf{poly}(n)]\subsetneq\mathsf{pCoT}[\mathsf{poly}(n)].$
@@ 5 Experiments
In this section, we provide empirical validation of our theoretical results on tasks with well-characterized complexity. Specifically, we study parallelizable tasks to empirically validate the efficiency of latent thought predicted in Section 3, and approximate counting and sampling tasks to demonstrate the effectiveness of CoT as shown in Section 4.
@@ 5.1 Experimental Setting
Fundamental algorithmic reasoning tasks.
@@
We use four problems. (1) Word problems for finite non-solvable groups: given a sequence of generators, the task is to evaluate their composition, which is $\mathsf{NC}^{1}$-complete (Barrington, 1986), also studied for Looped TF (Merrill and Sabharwal, 2025a). (2) sstt connectivity (STCON): given a directed graph $G=(V,E)$ and two vertices $s,t\in V$, the task is to decide whether $t$ is reachable from $s$, which belongs to $\mathsf{TC}^{1}$ (Gibbons and Rytter, 1989). (3) Arithmetic expression evaluation: given a formula consisting of $+,\times,-,/$ operations on integers, the task is to evaluate it. This problem is $\mathsf{TC}^{0}$-reducible to Boolean formula evaluation (Feng et al., 2023), which is $\mathsf{NC}^{1}$-complete (Buss, 1987). (4) Edit distance: given two strings $x$ and $y$, the task is to compute the minimum cost to transform $x$ into $y$. By reducing the dynamic programming formulation to shortest paths, this problem is in $\mathsf{TC}^{1}$ (Apostolico et al., 1990).
Approximate counting tasks.
We consider DNF counting and uniform sampling of graph colorings, both of which admit fully polynomial randomized approximation schemes for counting and sampling (FPRAS and FPAUS). Specifically, DNF counting admits an FPRAS via Monte Carlo sampling (Karp et al., 1989), while approximate counting and sampling of graph colorings admit an FPAUS based on rapidly mixing Markov chain Monte Carlo under suitable degree and color constraints (Jerrum, 1995).
Training strategy.
Since our primary objective is to study expressive power, we allow flexibility in optimization and training strategies. For CoT models, training is performed with supervision from explicit sequential algorithms. For fewer CoT steps, we compare two strategies: uniformly selecting steps from the indices of the complete trajectory (Bavandpour et al., 2025), and stepwise internalization (distillation) methods (Deng et al., 2024). For latent thought, we observe that looped TF is easier to train than Coconut, and therefore adopt looped TF as our instantiation of latent thought, with curriculum learning applied to certain tasks.
Table 2: Accuracy (%) of CoT and looped TF on parallelizable tasks across different numbers of iterations. Here, $n$ denotes the problem size. For CoT, we report the best accuracy achieved across the two training strategies.
Task $\bm{n}$ Looped Transformer Chain of Thought
2 4 6 8 8 16 32 64
Word Problem 64 0.8 0.8 100.0 100.0 0.8 0.8 100.0 100.0
Graph Connectivity 32 80.8 95.8 99.0 99.0 81.0 81.4 88.2 100.0
Arithmetic Evaluation 32/16 43.7 99.4 99.5 99.7 47.3 47.6 48.2 82.5
Edit Distance 32/16 57.3 72.9 86.2 90.7 76.5 80.9 87.5 94.8
@@ 5.2 Results
Table 2 reports results on parallelizable tasks, comparing latent thought and CoT under varying numbers of iterations. Latent thought solves the problems with fewer iterations than CoT requires to reach comparable performance. These empirical results are consistent with our theoretical analysis: latent thought supports efficient parallel reasoning, in contrast to the inherently sequential nature of CoT.
Figure 6: Accuracy of looped TFs on the arithmetic evaluation (top) and the connectivity (bottom). Each curve shows the performance for a fixed loop count $r$ as the input size $n$ increases.
We also evaluate the relationship between performance, input size, and the number of iterations, as in prior studies (Sanford et al., 2024b; Merrill and Sabharwal, 2025a). Figure 6 presents our results for looped TFs, illustrating that as the input size $n$ increases, the number of loops required to maintain high accuracy grows only logarithmically, supporting our theoretical claim in the (poly-)logarithmic regime.
Figure 7: Performance of Looped TF and CoT. Top: Relative error for counting (left) and sampling (right). Bottom: Empirical distributions for approximate sampling of graph colorings.
Figure 7 shows the results on the approximate counting and approximate sampling tasks. For approximate counting, CoT performs Monte Carlo estimation: the effective number of samples is given by the product of the number of reasoning steps per trial and the number of independent trials. Accordingly, increasing the iteration count corresponds to increasing this total sample budget. For approximate sampling, CoT emulates Markov chain Monte Carlo, where the iteration axis represents the number of Markov chain steps per trial. In contrast, for looped TF on both tasks, the iteration count simply corresponds to the number of loop iterations applied within a single trial. These results support the separation advantages of CoT with stochastic decoding.
@@ 6 Conclusion
We formally analyze the computational capabilities of chain-of-thought and latent thought reasoning, providing a rigorous comparison that reveals their respective strengths and limitations. Specifically, we show that latent thought enables efficient parallel computation, whereas CoT enables randomized approximate counting. Our theoretical results are further supported by experiments. These insights offer practical guidance for reasoning model design. For future work, an important direction is to investigate whether techniques such as distillation can reduce the number of iterations without compromising computational power. Another promising avenue is to extend our analysis to diffusion language models, which possess both parallelizability and stochasticity. Moreover, extending the analysis to realistic downstream tasks remains an important direction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Formal Definitions
A.1 Notation
@@
A.2 Transformer Block
We define the computational components of a Transformer block using the notation of (Merrill and Sabharwal, 2025a). Let $\mathbb{F}_{s}$ denote the set of $s$-bit floating-point numbers with truncated arithmetic (Definition B.1).
Definition A.1 (Transformer).
@@ A Transformer consists of the following components:
    A multi-head self-attention layer $\mathbf{SA}:\mathbb{F}_{s}^{m\times N}\to\mathbb{F}_{s}^{m\times N}$ for arbitrary sequence length $N$, parameterized by a matrix ${\bm{O}}:\mathbb{F}_{s}^{s\times H}\to\mathbb{F}_{s}^{m}$ and, for each head $h\in[H]$ with head size $s$, matrices ${\bm{Q}}_{h},{\bm{K}}_{h},{\bm{V}}_{h}:\mathbb{F}_{s}^{m}\to\mathbb{F}_{s}^{s}$. Given an input ${\bm{x}}_{i}\in\mathbb{F}_{s}^{m}$ for each position $i\in[N]$, it computes the query $\mathbf{q}_{i,h}={\bm{Q}}_{h}{\bm{x}}_{i}$, key ${\bm{k}}_{i,h}={\bm{K}}_{h}{\bm{x}}_{i}$, and value $\mathbf{v}_{i,h}={\bm{V}}_{h}{\bm{x}}_{i}$, and outputs ${\bm{O}}\cdot({\bm{a}}_{i,1},\ldots,{\bm{a}}_{i,H}),$ where each attention output ${\bm{a}}_{i,h}$ is defined, for softmax function, as:
    ${\bm{a}}_{i,h}=\sum_{j=1}^{c(i)}\frac{\exp(\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j,h})}{Z_{i,h}}\cdot\mathbf{v}_{j,h},\quad Z_{i,h}=\sum_{j=1}^{c(i)}\exp(\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j,h}),$ (1)
    with $c(i)=i$ for causal attention and $c(i)=N$ for full attention. For the saturated hardmax attention (Merrill et al., 2022), each attention output ${\bm{a}}_{i,h}$ is defined as:
    ${\bm{a}}_{i,h}=\sum_{j\in M_{i,h}}\frac{1}{|M_{i,h}|}\,\mathbf{v}_{j,h},\quad M_{i,h}=\left\{j\in[c(i)]\;\middle|\;\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j,h}=\max_{j^{\prime}}\mathbf{q}_{i,h}^{\top}{\bm{k}}_{j^{\prime},h}\right\}.$ (2)
@@
A.3 Chain of Thought
Definition A.2 (CoT).
@@ A.3 Chain of Thought
$\mathrm{TF}_{\mathrm{dec}}\coloneq\mathbf{OUT}\circ(\operatorname{id}+\mathbf{FF}_{L})\circ(\operatorname{id}+\mathbf{SA}_{L})\circ\cdots\circ(\operatorname{id}+\mathbf{FF}_{1})\circ(\operatorname{id}+\mathbf{SA}_{1})\circ(\mathbf{WE}+\mathbf{PE}),$ (3)
@@
$f_{\mathrm{cot}}^{k+1}(x)\coloneq f_{\mathrm{cot}}^{k}(x)\cdot\mathrm{Dec}\,(\mathrm{TF}_{\mathrm{dec}}(f_{\mathrm{cot}}^{k}(x))),$ (4)
where $\cdot$ denotes concatenation, and $\mathrm{Dec}(\cdot)$ is a decoding function that maps the output logits to a token in $\mathcal{V}$}: in the deterministic model, $\mathrm{Dec}(z)\coloneq\arg\max_{i\in[|{\mathcal{V}}|]}z_{i}$; in the stochastic model, $\mathrm{Dec}(z)\sim\mathrm{Multinomial}({z_{i}}/{\sum_{j}z_{j}})$, assuming $z_{i}>0$ for all $i$. The final output of the CoT model after $T(n)$ steps is defined as the last output length $m$ tokens of $f_{\mathrm{cot}}^{T(n)}(x)$.
A.4 Continuous Thought
Definition A.3 (Coconut).